OpenStudy (anonymous):

multiply: ( 5 - 4i) ^2

OpenStudy (tkhunny):

Have you considered performing the indicated multiplication?

OpenStudy (mathmate):

Recall: i*i=\(i^2\)=-1 \[(5-4i)^2=5^2-40i+16i^2=?\]